OpenStudy (anonymous):

3x=4 x= last time plz show and explain thxs

OpenStudy (anonymous):

3x=4 x=4/3

OpenStudy (anonymous):

why is it 4/3?

OpenStudy (anonymous):

3x=4 divide both sides bye 1/3 (3x)/3=4/3 then x=4/3

OpenStudy (anonymous):

el 3 que multiplica a x se pasa a dividir al 4